# Some Results on Finitely Splitting Subtrees of Aronszajn Trees

### v1

@@ Abstract.
For any $2\leq n<\omega$, we introduce a forcing poset using generalized promises which adds a normal $n$-splitting subtree to a $(\geq\!n)$-splitting normal Aronszajn tree. Using this forcing poset, we prove several consistency results concerning finitely splitting subtrees of Aronszajn trees. For example, it is consistent that there exists an infinitely splitting Suslin tree whose topological square is not Lindelöf, which solves an open problem due to Marun. For any $2<n<\omega$, it is consistent that every $(\geq\!n)$-splitting normal Aronszajn tree contains a normal $n$-splitting subtree, but there exists a normal infinitely splitting Aronszajn tree which contains no $(<\!n)$-splitting subtree. To show the latter consistency result, we prove a forcing iteration preservation theorem related to not adding new small-splitting subtrees of Aronszajn trees.
@@
The author acknowledges support from the Simons Foundation under the Travel Support for Mathematicians gift 631279
@@ 1. Introduction
In this article we study finitely splitting subtrees of Aronszajn trees. We prove consistency results related to the issue of the existence or non-existence of uncountable downwards closed subtrees of Aronszajn trees whose elements all have some prescribed number of immediate successors. This topic has deep connections with fundamental notions in the theory of trees, such as the property of being Suslin, as well as having significance to topology. For example, recently Marun [Mar] has shown that an $\omega_{1}$-tree, equipped with a natural topology called the fine wedge topology, is Lindelöf if and only it does not contain an uncountable downwards closed finitely splitting subtree.
We introduce a forcing poset for adding an uncountable downwards closed normal subtree of a given Aronszajn tree with some specified arity of its elements using the technique of promises. The idea of a promise was introduced by Shelah [She98, Chapter V, Section 6] and incorporated in his forcing for specializing an Aronzajn tree without adding reals. Roughly speaking, a promise is an uncountable subtree of a derived tree of an Aronszajn tree which can be used in a forcing notion to put constraints on the growth of the working part of a condition. While introduced explicitly by Shelah, a version of promises appeared earlier within Jensen’s proof of the consistency of Suslin’s hypothesis with the continuum hypothesis ([DJ74, page 108]). More recently, the use of promises to force subtrees of Aronszajn trees appears in work of Moore and Lamei Ramandi ([Moo07], [LRM18]). Our forcing for adding a subtree uses a generalized version of promises which is similar to that introduced by Abraham-Shelah [AS93, Definition 4.1] to specialize an Aronszajn tree while preserving Suslin trees.
@@
Let $T$ be an $\omega_{1}$-tree, which means a tree of height $\omega_{1}$ with countable levels. For any $0<n\leq\omega$, $T$ is $n$-splitting if every element of $T$ has exactly $n$-many immediate successors. Similar properties such as $(\leq\!n)$-splitting, $(<\!n)$-splitting, and $(\geq\!n)$-splitting (when $n$ is finite) have the obvious meaning. By a subtree of $T$ we mean any set $W\subseteq T$ considered as a tree with the order $<_{T}\cap\ W^{2}$. We will mostly, but not exclusively, be interested in subtrees which are downwards closed in the obvious sense. For any $x\in T$, let $T_{x}$ be the subtree $\{y\in T:x\leq_{T}y\}$.
For any $x\in T$, we write $\mathrm{ht}_{T}(x)$ for the height of $x$ in $T$. Let $T_{\alpha}$, or level $\alpha$ of $T$, denote the set of all elements of $T$ with height $\alpha$, for any $\alpha<\omega_{1}$. Let $T\upharpoonright A=\{x\in T:\mathrm{ht}_{T}(x)\in A\}$, for any $A\subseteq\omega_{1}$. If $b$ is a cofinal branch of $T$ and $\alpha<\omega_{1}$, let $b(\alpha)$ be the unique element of $b\cap T_{\alpha}$. Whenever $x\in T$ and $\beta<\mathrm{ht}_{T}(x)$, $x\upharpoonright\beta$ is the unique element below $x$ with height $\beta$. Similarly, $X\upharpoonright\beta=\{x\upharpoonright\beta:x\in X\}$ and $\vec{x}\upharpoonright\beta=(x_{0}\upharpoonright\beta,\ldots,x_{n-1}\upharpoonright\beta)$ whenever $X\subseteq T_{\alpha}$ and $\vec{x}=(x_{0},\ldots,x_{n-1})\in T_{\alpha}^{n}$ for some $\beta<\alpha<\omega_{1}$. For any $x\in T$, we write $\mathrm{Imm}_{T}(x)$ for the set of immediate successors of $x$. We say that $T$ is normal if it has a root, is Hausdorff, every element has incomparable elements above it, and every element has uncountably many elements above it (if $T$ is Aronszajn, then the third assumption is subsumed by the fourth).
@@ 2. Finitely Splitting Subtrees and Lindelöf Trees
Start with a special Aronszajn tree $T$ which is normal and infinitely splitting. We build a normal $\omega_{1}$-tree $U$ satisfying that $U\upharpoonright(\text{Lim}(\omega_{1})\cup\{0\})=T$. For any $x\in T$, insert $\omega$-many new levels in between $x$ and its immediate successors in $T$ which is a copy of the tree $({}^{<\omega}n\setminus\{\emptyset\},\subset)$. Now pick distinct cofinal branches $\langle b_{n}:n<\omega\rangle$ of this copy of ${}^{<\omega}n\setminus\{\emptyset\}$ so that every element of the copy belongs to some branch. Enumerating the immediate successors of $x$ in $T$ as $\langle y_{n}:n<\omega\rangle$, specify $y_{n}$ to be the unique upper bound of $b_{n}$ for all $n<\omega$. It is routine to check that the tree $U$ obtained in this manner is normal and $n$-splitting. Since $U$ restricted to the club set of countable limit ordinals is special, by Lemma 1.1(c) so is $U$. ∎
@@
Begin by letting $U^{0}=T$. Suppose that $U^{\alpha}$ is defined for some fixed $\alpha<\omega_{1}$. Define $U^{\alpha+1}$ which extends $U^{\alpha}$ by adding above each $x\in U^{\alpha}_{\alpha}$ infinitely many disjoint trees, which we denote by $U(\alpha,x,n)$ for $n<\omega$, each of them isomorphic to $T$ minus the root of $T$. Since $T$ is special, we can fix a specializing function $\pi_{\alpha,x,n}:U(\alpha,x,n)\to\omega$ for each such $x$ and $n$. This completes the constriction of $U^{\alpha+1}$, which easily satisfies the required properties. Now assume that $\delta<\omega_{1}$ is a limit ordinal and $U^{\alpha}$ is defined for all $\alpha<\delta$. Let $U^{\delta}=\bigcup_{\alpha<\delta}U^{\alpha}$. Again, it is straightforward to verify that $U^{\delta}$ is as required. This completes the construction.
@@
To show that $U$ is special, we define a map $\pi_{U}:U\to{}^{<\omega}\omega\setminus\{\emptyset\}$ by induction as follows. For any $z\in U^{0}=T$, let $\pi_{U}(z)=\langle\pi(z)\rangle$. Now consider $z\in U\setminus U^{0}$ and assume that $\pi_{U}$ is defined for all $y\in U$ such that $\alpha_{y}<\alpha_{z}$. In particular, we have defined $\pi_{U}(x_{z})$. Let $\pi_{U}(z)$ be the concatenation $\pi_{U}(x_{z})^{\frown}\pi_{\beta_{z},x_{z},n_{z}}(z)$.
@@
$\pi_{U}(y)=\pi_{U}(x_{y})^{\frown}\pi_{\beta_{y},x_{y},n_{y}}(y)\neq\pi_{U}(x_{z})^{\frown}\pi_{\beta_{z},x_{z},n_{z}}(z)=\pi_{U}(z).$
@@
$\pi_{U}(y)=\pi_{U}(x_{y})^{\frown}\pi_{\beta_{y},x_{y},n_{y}}(y)\neq\pi_{U}(x_{z})^{\frown}\pi_{\beta_{z},x_{z},n_{z}}(z)=\pi_{U}(z).$
@@
Let $2\leq m<n<\omega$ and suppose that $T$ is a $(\geq\!n)$-splitting $\omega_{1}$-tree which contains a $(\leq\!m)$-splitting subtree. Then $T$ also contains a (non-normal) $n$-splitting subtree, which can be obtained by adding enough terminal elements above each member of the $(\leq\!m)$-splitting subtree. The triviality of this observation shows the importance of the assumption of normality on the subtrees.
@@
We now turn to topological applications of finitely splitting subtrees. Let $T$ be an $\omega_{1}$-tree. For any $x\in T$ let $x\!\!\uparrow\ =\{y\in T:x\leq_{T}y\}$, and for any set $X\subseteq T$ let $X\!\!\uparrow\ =\bigcup\{x\!\!\uparrow\ :x\in X\}$.
@@
Going forward we refer to an $\omega_{1}$-tree which is Lindelöf in the fine wedge topology as a Lindelöf tree. Note that if an $\omega_{1}$-tree $T$ is finitely splitting, then the fine wedge topology on $T$ is discrete and hence $T$ is non-Lindelöf. Also, if $T$ has a cofinal branch, then $T$ is non-Lindelöf ([Mar, Lemma 4.2]). Less trivial examples of non-Lindelöf trees are those which are Aronszajn and infinitely splitting. Marun observed that there exists a normal Aronszajn tree which is infinitely splitting and non-Lindelöf, and that any infinitely splitting Suslin tree is Lindelöf.
@@
There exists a normal infinitely splitting non-Lindelöf special Aronszan tree.
@@
In this article, we address the following questions of Marun ([Mar24], [Mar]).
@@ 3. A Preservation Theorem for Iterated Forcing
In this section we prove a preservation theorem for iterated forcing related to not adding uncountable small-splitting subtrees of a given $\omega_{1}$-tree. One of our main consistency results, proven in Section 11, is that for any $2<n<\omega$, it is consistent that every $(\geq\!n)$-splitting normal Aronszajn tree contains an uncountable downwards closed normal $n$-splitting subtree, but there exists an infinitely splitting normal Aronszajn tree which contains no uncountable downwards closed $(<\!n)$-splitting subtree. In Section 6 we introduce a forcing for adding such an $n$-splitting subtree, and in Section 7 we show that this forcing does not add a $(<\!n)$-splitting subtree to any $\omega_{1}$-tree which did not already contain one. But in order to obtain the above consistency result, we need to be able to iterate such forcings while preserving these properties. To show that this can be done is the purpose of the next theorem.
@@
Let $T$ be an $\omega_{1}$-tree and let $1<n\leq\omega$. Suppose that $T$ does not contain an uncountable downwards closed $(<\!n)$-splitting subtree. Then the property of a forcing poset that it is proper and forces that $T$ does not contain an uncountable downwards closed $(<\!n)$-splitting subtree is preserved by countable support forcing iterations.
@@
Then there exists some $q^{+}\in\mathbb{P}_{\delta}$ such that $q^{+}\upharpoonright\alpha=q$, $q^{+}$ is $(N,\mathbb{P}_{\delta})$-generic, and $q^{+}\Vdash_{\delta}\dot{p}\in\dot{G}_{\delta}$.
@@
Fix $T$ and $n$ as in the statement of the theorem. We prove by induction on $\beta$ that if $\langle\mathbb{P}_{i},\dot{\mathbb{Q}}_{j}:i\leq\beta,\ j<\beta\rangle$ is a countable support forcing iteration such that for all $i<\beta$, $\mathbb{P}_{i}$ forces that ˙$\dot{\mathbb{Q}}_{i}$ is proper and ˙$\dot{\mathbb{Q}}_{i}$ forces that $T$ does not contain an uncountable downwards closed $(<\!n)$-splitting subtree, then $\mathbb{P}_{\beta}$ is proper and forces that $T$ does not contain an uncountable downwards closed $(<\!n)$-splitting subtree. Without loss of generality we may assume that for all $\gamma\leq\beta$, $\mathbb{P}_{\gamma}$ is separative, for otherwise there exists an equivalent forcing iteration which has this property (this step is not necessary, but it eliminates the need for an additional lemma).
The cases in which $\beta$ is equal to 00 or is a successor ordinal are immediate. So assume that $\beta$ is a limit ordinal. By the inductive hypothesis, for all $\gamma<\beta$, $\mathbb{P}_{\gamma}$ is proper and forces that $T$ does not contain an uncountable downwards closed $(<\!n)$-splitting subtree.
Suppose that $p\in\mathbb{P}_{\beta}$ and $p$ forces that $\dot{U}$ is a downwards closed $(<\!n)$-splitting subtree of $T$. We find an extension of $p$ which forces that $\dot{U}$ is countable.
@@
Clearly, $W$ is a downwards closed subtree of $T$ and $u\Vdash_{\mathbb{P}_{\beta}/G_{\alpha}}W\subseteq\dot{U}$. Since $u\leq_{\beta}p$ and $p$ forces that $\dot{U}$ is $(<\!n)$-splitting, $W$ must be $(<\!n)$-splitting.
By the inductive hypothesis, $T$ does not contain an uncountable downwards closed $(<\!n)$-splitting subtree in $V[G_{\alpha}]$. Therefore, $W$ must be countable. Fix some $\gamma<\omega_{1}$ such that $W\subseteq T\upharpoonright\gamma$. Now consider any $x\in T_{\gamma}$. Then $x\notin W$, so by the definition of $W$ there exists some $p_{x}\leq_{\beta}u$ in $\mathbb{P}_{\beta}/G_{\alpha}$ such that $p_{x}\Vdash_{\mathbb{P}_{\beta}/G_{\alpha}}x\notin\dot{U}$. Since $p_{x}\in\mathbb{P}_{\beta}/G_{\alpha}$, $p_{x}\upharpoonright\alpha\in G_{\alpha}$. Find $v_{x}\in G_{\alpha}$ which forces in $\mathbb{P}_{\alpha}$ over $V$ that $p_{x}\in\mathbb{P}_{\beta}/\dot{G}_{\alpha}$ and $p_{x}\Vdash_{\mathbb{P}_{\beta}/\dot{G}_{\alpha}}^{V[\dot{G}_{\alpha}]}x\notin\dot{U}$. By separativity, $v_{x}\leq_{\alpha}p_{x}\upharpoonright\alpha$. Now let $q_{x}=v_{x}\cup p_{x}\upharpoonright[\alpha,\beta)$. Then $q_{x}\leq_{\beta}p_{x}\leq_{\beta}u$. It is easy to check that $q_{x}\Vdash^{V}_{\beta}x\notin\dot{U}$. As $u\in D$ and $D$ is open, $q_{x}\in D$. $\dashv_{\ \text{Claim}}$
@@
This completes the construction. Let $q=\bigcup_{m}q_{m}$, which is in $\mathbb{P}_{\beta}$. We argue that $q\leq_{\beta}p$, $q$ is $(N,\mathbb{P}_{\beta})$-generic, and $q$ forces that $\dot{U}$ is countable. Fix a generic filter $G_{\beta}$ for $\mathbb{P}_{\beta}$ which contains $q$ and let $U=\dot{U}^{G_{\beta}}$. For each $m<\omega$ let $G_{\alpha_{m}}=G_{\beta}\upharpoonright\alpha_{m}$ and $p_{m}=\dot{p}_{m}^{G_{\alpha_{m}}}$.

### v2

@@ Abstract.
For any $2\leq n<\omega$, we introduce a forcing poset using generalized promises which adds a normal $n$-splitting subtree to a $(\geq\!n)$-splitting normal Aronszajn tree. Using this forcing poset, we prove several consistency results concerning finitely splitting subtrees of Aronszajn trees. For example, it is consistent that there exists an infinitely splitting Suslin tree whose topological square is not Lindelöf, which solves an open problem due to Marun. For any $2<n<\omega$, it is consistent that every $(\geq\!n)$-splitting normal Aronszajn tree contains a normal $n$-splitting subtree, but there exists a normal infinitely splitting Aronszajn tree which contains no (<n)(<\!n)-splitting subtree. To show the latter consistency result, we prove a forcing iteration preservation theorem related to not adding new small-splitting subtrees of Aronszajn trees.
@@ 1. Introduction
In this article we study finitely splitting subtrees of Aronszajn trees. We prove consistency results related to the issue of the existence or non-existence of uncountable downwards closed subtrees of Aronszajn trees whose elements all have some prescribed number of immediate successors. This topic has deep connections with fundamental notions in the theory of trees, such as the property of being Suslin, as well as having significance to topology. For example, recently Marun [Mar24b] has shown that an $\omega_{1}$-tree, equipped with a natural topology called the fine wedge topology, is Lindelöf if and only it does not contain an uncountable downwards closed finitely splitting subtree.
We introduce a forcing poset for adding an uncountable downwards closed normal subtree of a given Aronszajn tree with some specified arity of its elements using the technique of promises. The idea of a promise was introduced by Shelah [She98, Chapter V, Section 6] and incorporated in his forcing for specializing an Aronzajn tree without adding reals. Roughly speaking, a promise is an uncountable subtree of a derived tree of an Aronszajn tree which can be used in a forcing notion to put constraints on the growth of the working part of a condition. While introduced explicitly by Shelah, a version of promises appeared earlier within Jensen’s proof of the consistency of Suslin’s hypothesis with the continuum hypothesis ([DJ74, page 108]). More recently, the use of promises to force subtrees of Aronszajn trees appears in work of Lamei Ramandi and Moore ([Moo07], [LRM18]). Our forcing for adding a subtree uses a generalized version of promises which is similar to that introduced by Abraham-Shelah [AS93, Definition 4.1] to specialize an Aronszajn tree while preserving Suslin trees.
@@
Let $T$ be an $\omega_{1}$-tree, which means a tree of height $\omega_{1}$ with countable levels. For any $0<n\leq\omega$, $T$ is $n$-splitting if every element of $T$ has exactly $n$-many immediate successors. Similar properties such as $(\leq\!n)$-splitting, (<n)(<\!n)-splitting, and $(\geq\!n)$-splitting (when $n$ is finite) have the obvious meaning. By a subtree of $T$ we mean any set $W\subseteq T$ considered as a tree with the order $_{T}\cap\ W^{2}$. We will mostly, but not exclusively, be interested in subtrees which are downwards closed in the obvious sense. For any $x\in T$, let $T_{x}$ be the subtree $\{y\in T:x\leq_{T}y\}$.
For any $x\in T$, we write $\mathrm{ht}_{T}(x)$ for the height of $x$ in $T$. Let $T_{\alpha}$, or level $\alpha$ of $T$, denote the set of all elements of $T$ with height $\alpha$, for any $\alpha<\omega_{1}$. Let $T\upharpoonright A=\{x\in T:\mathrm{ht}_{T}(x)\in A\}$, for any $A\subseteq\omega_{1}$. If $b$ is a cofinal branch of $T$ and $\alpha<\omega_{1}$, let $b(\alpha)$ be the unique element of $b\cap T_{\alpha}$. Whenever $x\in T$ and $\beta<\mathrm{ht}_{T}(x)$, $x\upharpoonright\beta$ is the unique element below $x$ with height $\beta$. Similarly, $X\upharpoonright\beta=\{x\upharpoonright\beta:x\in X\}$ and $\vec{x}\upharpoonright\beta=(x_{0}\upharpoonright\beta,\ldots,x_{n-1}\upharpoonright\beta)$ whenever $X\subseteq T_{\alpha}$ and $\vec{x}=(x_{0},\ldots,x_{n-1})\in T_{\alpha}^{n}$ for some $\beta<\alpha<\omega_{1}$.
For any $x\in T$, we write $\mathrm{Imm}_{T}(x)$ for the set of immediate successors of $x$. We say that $T$ is normal if it has a root, is Hausdorff, every element has incomparable elements above it, and every element has uncountably many elements above it (if $T$ is Aronszajn, then the third assumption is subsumed by the fourth). A subset of $T$ is dense open if it is dense open in the forcing poset consisting of $T$ partially ordered by the reverse of the tree ordering.
@@ 2. Finitely Splitting Subtrees and Lindelöf Trees
Start with a special Aronszajn tree $T$ which is normal and infinitely splitting. We build a normal $\omega_{1}$-tree $U$ satisfying that $U\upharpoonright(\text{Lim}(\omega_{1})\cup\{0\})=T$. For any $x\in T$, insert $\omega$-many new levels in between $x$ and its immediate successors in $T$ which is a copy of the tree $({}^{<\omega}n\setminus\{\emptyset\},\subset)$. Now pick distinct cofinal branches $\langle b_{k}:k<\omega\rangle$ of this copy of $^{<\omega}n\setminus\{\emptyset\}$ so that every element of the copy belongs to some branch. Enumerating the immediate successors of $x$ in $T$ as $\langle y_{k}:k<\omega\rangle$, specify $y_{k}$ to be the unique upper bound of $b_{k}$ for all $k<\omega$. It is routine to check that the tree $U$ obtained in this manner is normal and $n$-splitting. Since $U$ restricted to the club set of countable limit ordinals is special, by Lemma 1.1(c) so is $U$. ∎
@@
Begin by letting $U^{0}=T$. Suppose that $U^{\alpha}$ is defined for some fixed $\alpha<\omega_{1}$. Define $U^{\alpha+1}$ which extends $U^{\alpha}$ by adding above each $x\in U^{\alpha}_{\alpha}$ infinitely many pairwise disjoint trees, which we denote by $U(\alpha,x,n)$ for $n<\omega$, each of them isomorphic to $T$ minus the root of $T$. Since $T$ is special, we can fix a specializing function $\pi_{\alpha,x,n}:U(\alpha,x,n)\to\omega$ for each such $x$ and $n$. This completes the constriction of $U^{\alpha+1}$, which easily satisfies the required properties. Now assume that $\delta<\omega_{1}$ is a limit ordinal and $U^{\alpha}$ is defined for all $\alpha<\delta$. Let $U^{\delta}=\bigcup_{\alpha<\delta}U^{\alpha}$. Again, it is straightforward to verify that $U^{\delta}$ is as required. This completes the construction.
@@
To show that $U$ is special, we define a map $\pi_{U}:U\to{}^{<\omega}\omega\setminus\{\emptyset\}$ by induction as follows. For any $z\in U^{0}=T$, let $\pi_{U}(z)=\langle\pi(z)\rangle$. Now consider $z\in U\setminus U^{0}$ and assume that $\pi_{U}(y)$ is defined for all $y\in U$ such that $\alpha_{y}<\alpha_{z}$. In particular, we have defined $\pi_{U}(x_{z})$. Let $\pi_{U}(z)$ be the concatenation $\pi_{U}(x_{z})^{\frown}\langle\pi_{\beta_{z},x_{z},n_{z}}(z)\rangle$.
@@
$\pi_{U}(y)=\pi_{U}(x_{y})^{\frown}\langle\pi_{\beta_{y},x_{y},n_{y}}(y)\rangle\neq\pi_{U}(x_{z})^{\frown}\langle\pi_{\beta_{z},x_{z},n_{z}}(z)\rangle=\pi_{U}(z).$
@@
$\pi_{U}(y)=\pi_{U}(x_{y})^{\frown}\langle\pi_{\beta_{y},x_{y},n_{y}}(y)\rangle\neq\pi_{U}(x_{z})^{\frown}\langle\pi_{\beta_{z},x_{z},n_{z}}(z)\rangle=\pi_{U}(z).$
@@
Let $2\leq m<n<\omega$ and suppose that $T$ is a $(\geq\!n)$-splitting $\omega_{1}$-tree which contains a $(\leq\!m)$-splitting subtree. Then $T$ also contains a (non-normal) $n$-splitting subtree, which can be obtained by adding new elements with successor height to the $(\leq\!m)$-splitting subtree to ensure $n$-splitting, but at limit levels only including members of the $(\leq\!m)$-splitting subtree. The triviality of this observation displays the importance of the assumption of normality on the subtrees.
@@
We now turn to topological applications of finitely splitting subtrees. Let $T$ be an $\omega_{1}$-tree. For any $x\in T$ let $x{\uparrow}$ denote the set $\{y\in T:x\leq_{T}y\}$, and for any $X\subseteq T$ let $X{\uparrow}$ denote the union $\bigcup\{x{\uparrow}:x\in X\}$.
@@
Going forward we refer to an $\omega_{1}$-tree which is Lindelöf in the fine wedge topology as a Lindelöf tree. Note that if an $\omega_{1}$-tree $T$ is finitely splitting, then the fine wedge topology on $T$ is discrete and hence $T$ is non-Lindelöf. Also, if $T$ has a cofinal branch, then $T$ is non-Lindelöf ([Mar24b, Lemma 4.2]). Less trivial examples of non-Lindelöf trees are those which are Aronszajn and infinitely splitting. Marun observed that there exists a normal Aronszajn tree which is infinitely splitting and non-Lindelöf, and that any infinitely splitting Suslin tree is Lindelöf.
@@
There exists a normal infinitely splitting non-Lindelöf special Aronszajn tree.
@@
In this article, we address the following questions of Marun ([Mar24a], [Mar24b]).
@@ 3. A Preservation Theorem for Iterated Forcing
In this section we prove a preservation theorem for iterated forcing related to not adding uncountable small-splitting subtrees of a given $\omega_{1}$-tree. One of our main consistency results, proven in Section 11, is that for any $2<n<\omega$, it is consistent that every $(\geq\!n)$-splitting normal Aronszajn tree contains an uncountable downwards closed normal $n$-splitting subtree, but there exists an infinitely splitting normal Aronszajn tree which contains no uncountable downwards closed (<n)(<\!n)-splitting subtree. In Section 6 we introduce a forcing for adding such an $n$-splitting subtree, and in Section 7 we show that this forcing does not add a (<n)(<\!n)-splitting subtree to any $\omega_{1}$-tree which did not already contain one. But in order to obtain the above consistency result, we need to be able to iterate such forcings while preserving these properties. To show that this can be done is the purpose of the next theorem.
@@
Let $T$ be an $\omega_{1}$-tree and let $1<n\leq\omega$. Assume that $T$ does not contain an uncountable downwards closed (<n)(<\!n)-splitting subtree. Suppose that $\langle\mathbb{P}_{i},\dot{\mathbb{Q}}_{j}:i\leq\beta,\ j<\beta\rangle$ is a countable support forcing iteration such that for all $\beta, i\mathbb{P}_{i}$ forces that $\dot{\mathbb{Q}}_{i}$ is proper and $\dot{\mathbb{Q}}_{i}$ forces that $T$ does not contain an uncountable downwards closed (<n)(<\!n)-splitting subtree. Then $\mathbb{P}_{\beta}$ is proper and forces that $T$ does not contain an uncountable downwards closed (<n)(<\!n)-splitting subtree.
@@
Then there exists some $q^{+}\in\mathbb{P}_{\delta}$ such that $^{+}\upharpoonright\alpha=q, q+q^{+}$ is $(N,\mathbb{P}_{\delta})$-generic, and $q^{+}\Vdash_{\delta}\dot{p}\in\dot{G}_{\delta}$.
@@
We prove the statement by induction on $\beta$. Fix $T$, $n$, and $\langle\mathbb{P}_{i},\dot{\mathbb{Q}}_{j}:i\leq\beta,\ j<\beta\rangle$ as in the theorem. Without loss of generality we may assume that for all $\gamma\leq\beta$, $\mathbb{P}_{\gamma}$ is separative, for otherwise there exists an equivalent forcing iteration which has this property (this step is not necessary, but it eliminates the need for an additional lemma).
The cases in which $\beta$ is equal to 0 or is a successor ordinal are immediate. So assume that $\beta$ is a limit ordinal. By the inductive hypothesis, for all $\gamma<\beta$, $\mathbb{P}_{\gamma}$ is proper and forces that $T$ does not contain an uncountable downwards closed (<n)(<\!n)-splitting subtree.
Suppose that $p\in\mathbb{P}_{\beta}$ and $p$ forces that $\dot{U}$ is a downwards closed (<n)(<\!n)-splitting subtree of $T$. We find an extension of $p$ which forces that $\dot{U}$ is countable.
@@
Clearly, $W$ is a downwards closed subtree of $T$ and $u\Vdash_{\mathbb{P}_{\beta}/G_{\alpha}}W\subseteq\dot{U}$. Since $u\leq_{\beta}p$ and $p$ forces that $\dot{U}$ is (<n)(<\!n)-splitting, $W$ must be (<n)(<\!n)-splitting.
By the inductive hypothesis, $T$ does not contain an uncountable downwards closed (<n)(<\!n)-splitting subtree in $V[G_{\alpha}]$. Therefore, $W$ must be countable. Fix some $\gamma<\omega_{1}$ such that $W\subseteq T\upharpoonright\gamma$. Now consider any $x\in T_{\gamma}$. Then $x\notin W$, so by the definition of $W$ there exists some $p_{x}\leq_{\beta}u$ in $\mathbb{P}_{\beta}/G_{\alpha}$ such that $p_{x}\Vdash_{\mathbb{P}_{\beta}/G_{\alpha}}x\notin\dot{U}$. Since $p_{x}\in\mathbb{P}_{\beta}/G_{\alpha}$, $p_{x}\upharpoonright\alpha\in G_{\alpha}$. Find $v_{x}\in G_{\alpha}$ which forces in $\mathbb{P}_{\alpha}$ over $V$ that $p_{x}\in\mathbb{P}_{\beta}/\dot{G}_{\alpha}$ and $p_{x}\Vdash_{\mathbb{P}_{\beta}/\dot{G}_{\alpha}}^{V[\dot{G}_{\alpha}]}x\notin\dot{U}$. By separativity, $v_{x}\leq_{\alpha}p_{x}\upharpoonright\alpha$. Now let $q_{x}=v_{x}\cup p_{x}\upharpoonright[\alpha,\beta)$. Then $q_{x}\leq_{\beta}p_{x}\leq_{\beta}u$. It is easy to check that $q_{x}\Vdash^{V}_{\beta}x\notin\dot{U}$. As $u\in D$ and $D$ is open, $q_{x}\in D$. $\dashv_{\ \text{Claim}}$
@@
This completes the construction. Let $q=\bigcup_{m}q_{m}$, which is in $\mathbb{P}_{\beta}$. We argue that $\leq_{\beta}p, qq$ is $(N,\mathbb{P}_{\beta})$-generic, and $q$ forces that $\dot{U}$ is countable. Fix a generic filter $G_{\beta}$ for $\mathbb{P}_{\beta}$ which contains $q$ and let $U=\dot{U}^{G_{\beta}}$. For each $m<\omega$ let $G_{\alpha_{m}}=G_{\beta}\upharpoonright\alpha_{m}$ and $p_{m}=\dot{p}_{m}^{G_{\alpha_{m}}}$.